A Bibliography of Supercomputing ’95

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

12 August 2014 Version 1.05

Abstract

This bibliography records articles presented at the Supercomputing ’95 conference.

Title word cross-reference

N [BH95].

*Really* [Wul95].

-Body [BH95].

1000 [SSM^{+}95]. 12 [MFC^{+}95]. 1990s [Goo95].

3D [BMWY95].

94 [Por95].

Access [BDL95, NMkL^{+}95, Rin95]. ACM [Kar95]. ACM/IEEE [Kar95]. Activities [Wra95]. Adaptable [KMQ95]. Adaptive [KB95]. Address [ENSO95]. Admission [FM95]. Advanced [Iwa95, Sch95]. Alamos [SD95]. Algorithm [HL95, LKC95]. Algorithms [FAJL^{+}95, GK95]. America [Hod95]. Analysis [AMCA^{+}95, BA95, GLL95, KK95a]. Anisotropies [BB95]. Anxiety [Mil95]. Application [AFMS95, CS95, HS95b, SS95]. Applications [AS95, CACR95, ENSO95, FM95, HS95c, LM^{+}95, Zie95a]. Applications-Driven [ENSO95]. Approach [Ayg95, Gar95, HS95a, Lab95, WR95, WS95]. Approaches [NMkL^{+}95]. Approximate [GKS95]. Archetypes [VG95]. Architectural [JHP95, RSS^{+}95]. Architecture [KPG95, KMQ95, HS95c]. Architecture-Adaptable [KMQ95].

Architectures [TRM^{+}95]. Area [SD95]. Array [DHSK95, GLL95]. Astrophysical [MT95]. ATD [Sch95]. ATM [HHM95]. Atmospheric [DS95, FAJL^{+}95]. Automata [Lil95]. Automated [BSM95]. Automatic [Ayg95, Gar95, KK95b, Lab95].

Back [GK95]. Balancing [BH95]. Bears [BASW95]. Behavior [HS95b]. Benefits [ENSO95]. Bisection [Bar95]. Body [BH95, MT95]. Boundary [NK95, WSL95]. Bursty [FM95].

CA [Kar95]. Call [FM95]. Careers [Ano95]. Case [KMQ95, NK95]. Cellular [Lil95]. Center [Kar95, Por95]. CFD [KPG95]. Challenge [BMWY95]. Challenges [Iwa95]. Characteristics [CACR95]. Chemical [CS95, DS95]. Chromodynamics [YNF^{+}95]. Circulation [KMQ95]. Classroom [Lil95, MFC^{+}95]. Clothes [Pan95]. Clustering [ENSO95]. Clusters [tBC95, HHM95]. CMB [BB95]. Coarse [HAM^{+}95]. Coarse-Grain [HAM^{+}95]. Code [VG95, BMWY95]. Coherent [KSB95]. Collective [HHM95, SCJ^{+}95]. Commercial [SBW95]. Communication [HHM95, RSS^{+}95, Rin95]. Community [Zie95a, Zie95b]. Compilation [AMCA^{+}95, AS95]. Compiler [GMS^{+}95, HAM^{+}95, KPG95]. Compiling [TRM^{+}95]. Components [LKC95]. Computational [BDL95, CT95, HS95a, PJ95, Por95]. Computations [JHP95]. Computer [Ano95, Don95a, MT95, SSM^{+}95]. Computers [WS95]. Computing [Goo95, LM^{+}95, MPS95, NMkL^{+}95, NK95, Rin95, Wul95]. Concepts [Cur95]. Concurrent [KA95]. Conference [Kar95]. Confront [BASW95]. Connected [LKC95]. Connection [SB95]. Consistency [KSB95]. Control [FM95, Goo95]. Controlling [SS95]. Convention [Kar95]. Convex [SSM^{+}95]. Correction [FAJL^{+}95]. Correlated [FM95]. Cray [BMWY95]. Cup [Hod95]. Curricula [CT95]. Curriculum [Ste95].

Data [AMCA^{+}95, Ayg95, BH95, FAJL^{+}95, Gar95, KK95b, Lab95, MPP^{+}95, Rin95]. Dataflow [GLL95]. Dataset [NMkL^{+}95]. Debugging [AFMS95]. December [Kar95]. Decoupled [TRM^{+}95]. Demonstration [Sch95]. Detecting [HAM^{+}95]. Detection [CJN^{+}95]. Developing [CT95, WSL95]. Development [AFMS95, Ste95, WR95]. Diego [Kar95]. Direct [GK95]. Directed [SCJ^{+}95]. Distributed [AS95, BA95, BDF^{+}95, FM95, HS95c, LDCZ95, NK95, PKZC95]. Distributed-Memory [HS95c, NK95]. Distributing [CS95]. Distribution [Ayg95, Gar95, Lab95]. Does [Don95a]. DOT [TMRP95]. Driven [ENSO95]. Dynamics [CSWM95].

EarthVision [CT95]. Eddy [WSL95].

Effect [KPG95]. Efficient [FAJL^{+}95, KA95]. Element [KMQ95, NK95]. Elliptic [LF95]. Embedded [LM^{+}95]. Emperor [Pan95]. Engineering [Ano95, Iwa95]. Environment [AMCA^{+}95, PJ95]. Environmental [NMkL^{+}95]. Environments [MPS95]. EUROPORT [Wra95]. Evaluation [SSM^{+}95, WR95]. Evolutionary [SB95]. Examination [Goo95]. Exchange [BDF^{+}95]. Execution [PKZC95]. Experience [CT95]. Experiences [Cur95]. Explicit [RSS^{+}95]. Exploiting [BH95]. Export [Goo95].

Fast [LKC95, PLC95]. Findings [Por95]. Fine [PKZC95]. Fine-Grained [PKZC95]. Finite [KMQ95]. Fitting [WB95]. Flattening [DHSK95]. Flops [SBW95]. Flow [LF95]. FM [PLC95]. Folding [HS95a]. Force [Sch95]. Fortran [KK95b]. Forward [GK95]. Future [BASW95, MFC^{+}95].

Gallon [Don95a]. Gateway [SD95]. GCV [WB95]. General [BB95]. Geostrophic [BMWY95]. Get [Don95a]. Ghost [SB95]. Gigabit [CS95, HS95c, SD95]. Global [DS95, WS95]. Goldilocks [BASW95]. Grain [HAM^{+}95, SS95]. Grained [PKZC95]. Grand [BMWY95]. GRAPE [MT95]. GRAPE-4 [MT95]. Graph [KK95a]. Graphs [HL95]. Grid [BMWY95]. Guide [Mil95].

Handy [Iwa95]. Hardware [KSB95]. Hardware-Coherent [KSB95]. Hear [Pan95]. Hierarchical [GKS95]. High [ACG95, Goo95, LM^{+}95, MPS95, NMkL^{+}95, PLC95, Por95, WS95, KK95b]. High-Performance [ACG95, Goo95, MPS95, NMkL^{+}95, WS95]. HIPPI [SD95]. HIPPI-SONET [SD95]. HPC [Pan95, Ste95]. HPCC [BDF^{+}95]. HPF [GMS^{+}95, Smi95]. Hybrid [PKZC95].

I/O [CSWM95, HS95c, Poo95, SCJ^{+}95]. IBM [BMWY95, BDG^{+}95, GMS^{+}95]. II [Zie95b]. Illinois [PLC95]. Imaging [MPP^{+}95]. Implementation [BMWY95]. Implementations [tBC95]. Incompressible [LF95]. Incremental [WS95]. Index [DHSK95]. Information [BCSS95, BDL95, BDF^{+}95, Mil95, Rin95]. Infrastructure [KB95]. Initiative [Por95]. Input [CACR95]. Input/Output [CACR95]. Instrumentation [WR95]. Instruments [Sko95]. Integrated [AMCA^{+}95]. Intelligent [BDL95]. Interactions [TMRP95]. Interface [VG95]. Interprocedural [AS95, HAM^{+}95]. Investigation [ENSO95]. Invited [Cur95, Don95a, Goo95, Hod95, Iwa95, Mil95, Pan95, Sch95]. Irregular [AS95].

Issues [Poo95].

Joint [Sch95]. JTF [Sch95].

Kernel [LF95]. Knowledge [Mil95].

Knowledge-Value [Mil95].

Languages [KA95, PKZC95]. Large [AFMS95, HS95b, NMkL^{+}95, WSL95]. Large-Scale [NMkL^{+}95]. Lattice [BDG^{+}95]. Layer [WSL95]. Layout [KK95b]. Lazy [KSB95]. Level [HL95, TD95]. Limitations [CSWM95]. Linear [BB95, GK95]. Living [MFC^{+}95]. Loads [BH95]. Locality [BH95]. Location [KA95].

Machine [TD95, SB95]. Machine-Level [TD95]. Machines [AS95, HS95c, LKC95]. Management [BDF^{+}95]. Manual [KPG95]. Many [Don95a]. Massively [WS95]. Matching [ACG95]. Matrix [GKS95]. Matrix-Vector [GKS95]. Mayors [Don95b]. Mechanics [HS95a]. Mechanisms [RSS^{+}95]. Memory [AS95, CBEL95, HS95c, LDCZ95, NK95, PKZC95, RSS^{+}95, SSM^{+}95, HS95b]. Mesh [KB95]. Message [BSM95, LDCZ95]. Message-Passing [BSM95]. Messages [PLC95]. Messaging [PLC95]. Method [NK95]. Methods [BDL95, GKS95, KB95]. Microparallelism [ACG95]. Miles [Don95a]. Misuse [CJN^{+}95]. Mobile [TD95]. Model [DS95, FM95, PKZC95]. Modeling [LKC95]. Modelling [KMQ95]. Models [AFMS95]. Modularity [SB95]. Molecular [CSWM95, TMRP95]. MONSTER [SB95]. MPCs [HHM95]. Multi [BMWY95, HL95]. Multi-Grid [BMWY95]. Multi-Level [HL95]. Multicast [HHM95]. Multicomputer [NK95]. Multicomputers [PKZC95]. Multigrid [LF95]. Multilevel [Bar95, KK95a]. Multiprocessors [ENSO95, HS95b, KSB95, KPG95, RSS^{+}95]. Myrinet [PLC95].

N [MT95]. N-body [MT95]. National [BDF^{+}95]. Navier [JHP95]. Need [Pan95]. Network [CS95, SS95, Cur95]. Networking [SD95, Zie95a, Zie95b]. Networks [LDCZ95]. Neural [SB95]. No [Pan95]. Novel [Ayg95, Gar95, Lab95]. Numerical [AFMS95, Hod95]. NWT [YNF^{+}95].

O [CSWM95, HS95c, Poo95, SCJ^{+}95]. Object [Gan95a, KA95, Gan95b]. Object-Oriented [KA95]. Ocean [KMQ95]. Optimization [CS95]. Optimizations [KPG95, Rin95]. Optimizing [TRM^{+}95]. Oriented [Gan95a, Gan95b, KA95]. Output [CACR95].

Package [LF95]. Panda [SCJ^{+}95]. Panel [Don95b]. Parallel [AMCA^{+}95, BA95, Bar95, BSM95, BB95, tBC95, CSWM95, CACR95, DS95, Gan95a, Gan95b, GKS95, GK95, KB95, LF95, LKC95, MPP^{+}95, NK95, Rin95, SSM^{+}95, VG95, WS95]. Parallelism [HAM^{+}95]. Parallelization [GLL95]. Parallelizing [HAM^{+}95, JHP95, JY95]. Partitioning [HL95, KK95a]. Passing [BSM95, LDCZ95]. IEEE [Kar95]. Performance [AMCA^{+}95, ACG95, BMWY95, BSM95, DS95, Goo95, KK95b, KPG95, LKC95, LM^{+}95, MPS95, NMkL^{+}95, PLC95, Sko95, SSM^{+}95, WS95]. Perspective [Smi95]. Phylogeny [JY95]. Pittsburgh [Por95]. Places [Cur95]. Plans [Cur95]. Platforms [JHP95]. PMRSB [Bar95]. Policy [Goo95]. Power [tBC95]. POWERParallel [BDG^{+}95]. Predicting [HS95b]. Prediction [BSM95, Hod95]. Price [Sko95]. Privatization [GLL95]. Problem [tBC95, JY95]. Proceedings [Kar95]. Process [CS95]. Processing [MPP^{+}95]. Processor [BH95]. Product [GKS95]. Program [DHSK95, GLL95, VG95]. Programming [Gan95a, Gan95b, KA95, PJ95, TD95]. Programs [AMCA^{+}95, BSM95]. Protein [ACG95, HS95a]. Protocols [CBEL95]. Purpose [MT95].

QCD [BDG^{+}95]. Quantum [YNF^{+}95]. Quasi [BMWY95]. Quasi-Geostrophic [BMWY95].

Radar [MPP^{+}95]. Recursive [Bar95]. Relative [AFMS95]. Relativity [BB95]. Release [KSB95]. Remotely [FAJL^{+}95]. Report [Por95]. Repository [NMkL^{+}95]. Research [Poo95, SB95]. Resources [Ste95]. Retrograde [BA95]. Revolution [GSS95, Mil95]. Road [BCSS95]. Robots [TD95]. Ruin [BCSS95].

San [Kar95]. Say [Pan95]. Scalable

[BDG^{+}95, CACR95, Poo95, SSM^{+}95]. Scale [HS95b, NMkL^{+}95]. Scheduling [WS95]. School [Por95]. Science [Ano95, CT95, Por95]. Scientific [NK95, PJ95]. SCIRun [PJ95]. SDSC [Ste95]. SDSU [Ste95]. Sensed [FAJL^{+}95]. Server [SCJ^{+}95]. Server-Directed [SCJ^{+}95]. Shared [ENSO95, HS95b, LDCZ95, RSS^{+}95, SSM^{+}95]. Shared-memory [HS95b]. Simulating [Sko95]. Simulation [KMQ95, WSL95, YNF^{+}95]. Simulations [BH95, MT95]. Size [SS95]. Smoothing [WB95]. Software [BDF^{+}95, GSS95, KB95, MPS95]. Solution [GK95]. Solver [LF95]. Some [Wul95]. SONET [SD95]. Source [VG95]. SP2 [BMWY95, BDG^{+}95, GMS^{+}95]. Space [ENSO95]. Spaceborne [MPP^{+}95]. Sparse [GK95]. Spatially [WSL95]. Spatially-Developing [WSL95]. Speaker [Cur95, Don95a, Goo95, Hod95, Iwa95, Mil95, Pan95, Sch95]. Special [MT95]. Special-Purpose [MT95]. Spectral [Bar95]. Splines [WB95]. SPP [SSM^{+}95]. SPP-1000 [SSM^{+}95]. Stability [tBC95]. Statistical [HS95a]. Steering [PJ95]. Stokes [JHP95]. Storm [CBEL95]. Structured [KB95, WR95]. Study [KMQ95, NK95]. Substitution [GK95]. Super [Iwa95, Wul95]. Supercomputer [GSS95]. Supercomputers [WB95]. Supercomputing [BASW95, Kar95, Por95]. Superhighway [BCSS95]. Support [KA95]. Surface [WB95]. Surveying [TMRP95]. Survival [Mil95]. Symbolic [GLL95]. Synergetic [KPG95]. System [BA95, tBC95, CBEL95, MPS95, WR95]. Systems [BDG^{+}95, GK95, SB95].

T3D [BMWY95]. Task [Sch95]. Teach [TD95]. Technologies [Iwa95]. Technology [Sch95, Zie95b]. Tereflops [SBW95]. Textbook [MFC^{+}95]. Theoretical [SB95]. Three [BASW95]. Thriving [Mil95]. TM [CJN^{+}95]. Tool [CBEL95]. Tools [MPS95]. Topologies [HHM95]. Tracer [DS95]. Traffic [FM95]. Transformation [DHSK95]. Transient [tBC95]. Transparency [KA95].

Undergraduate [Ste95]. UNICORN [CJN^{+}95]. UNICOS [CJN^{+}95]. USA [Kar95]. Use [Lil95]. User [Smi95]. Users [Pan95]. Using [GKS95, HAM^{+}95, NMkL^{+}95, Rin95, WB95, KMQ95, Ste95].

Value [Mil95]. Variety [JHP95]. Vector [GKS95]. Vendors [Pan95]. Versus [LDCZ95]. Virtual [HHM95, Iwa95]. Visualizing [CBEL95].

Watch [CBEL95]. Weather [Hod95]. Web [VG95]. Where [GSS95]. Wide [SD95]. Wide-Area [SD95]. Wind [Sko95]. Women [Ano95]. Workshop [Ano95, Gan95b, MPS95, Poo95, Smi95, Wra95]. Workstations [tBC95, LDCZ95, PLC95, SS95]. WWW [NMkL^{+}95].

Years [Por95].

References

Alpern:1995:MHP

[ACG95] Bowen Alpern, Larry Carter, and Kang Su Gatlin. Microparallelism and high-performance protein matching. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                                   //www.supercomp.org/sc95/
                                                   proceedings/549_LCAR/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Abramson:1995:RDA

[AFMS95] David Abramson, Ian Foster, John Michalakes, and Rok Sosič. Relative debugging and its application to the development of large numerical models. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                                   //www.supercomp.org/sc95/
                                                   proceedings/599_IFOS/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Adve:1995:ICP

[AMCA^{+}95] Vikram S. Adve, John Mellor-Crummey, Mark Anderson, Ken Kennedy, Jhy-Chun Wang, and Daniel A. Reed. An integrated compilation and performance analysis environment for data parallel programs. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/528_VADV/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Anonymous:1995:WCW

[Ano95] Anonymous. Workshop: Careers for women in computer science and engineering. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/MISC/WORKSHOP.

HTM#Women. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Agrawal:1995:ICI

[AS95] Gagan Agrawal and Joel Saltz. Interprocedural compilation of irregular applications for distributed memory machines. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/405_AGR/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Ayguade:1995:NAT

[Ayg95] Eduard Ayguade. A novel approach towards automatic data distribution. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/437_

JGAR/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Bal:1995:PRA

[BA95] Henri Bal and Victor Allis. Parallel retrograde analysis on a distributed system. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/

proceedings/463_HBAL/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Barnard:1995:PPM

[Bar95] Stephen T. Barnard. PMRSB:

Parallel multilevel recursive spectral bisection. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/458_SBAR/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Borchers:1995:GTB

[BASW95] Robert Borchers (moderator), Tom Anderson (panellist), Burton Smith (panellist), and Steven Wallach (panellist). Goldilocks and the Three Bears confront the future of supercomputing. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/MISC/
PANELS.HTM#Goldi. These pro-

ceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Bode:1995:PLG

[BB95] Paul Bode and Edmund Bertschinger. Parallel linear general relativity and CMB anisotropies. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/423_BODE/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Bace:1995:ISR

[BCSS95] Becky Bace (moderator), Gary Christoph (panelist), Tsutomu Shimomura (panelist), and Gene Spafford (panelist). Information superhighway or road to ruin? In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/MISC/
                                        PANELS.HTM#Bace. These pro-

ceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Browne:1995:DIM

[BDF^{+}95] Shirley Browne, Jack Dongarra, Geoffrey C. Fox, Ken Hawick, Ken Kennedy, Rick Stevens, Robert Olson, and Tom Rowan. Distributed information management in the National HPCC Software Exchange. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/588_

BROW/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Bernard:1995:LQI

[BDG^{+}95] C. Bernard, C. DeTar, S. Gottlieb, U. M. Heller, J. Hetrick, N. Ishizuka, L. Karkkainen, S. R. Lantz, K. Rummukainen, R. Sugar, D. Toussaint, and M. Wingate. Lattice QCD on the IBM Scalable POWERParallel Systems SP2. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/411_

CDT/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Berry:1995:CMI

[BDL95] Michael W. Berry, Susan T. Dumais, and Todd A. Letsche. Computational methods for intelligent information access. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/473_MBER/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Banicescu:1995:BPL

[BH95] Ioana Banicescu and Susan Flynn Hummel. Balancing processor loads and exploiting data locality in N-body simulations. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/594_BHUM/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Baillie:1995:IPG

[BMWY95] Clive F. Baillie, James C.

McWilliams, Jeffrey B. Weiss, and Irad Yavneh. Implementation and performance of a grand challenge 3D quasi-geostrophic multi-grid code on the Cray T3D and IBM SP2. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/424_

CLIV/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Block:1995:APP

[BSM95] Robert J. Block, Sekhar Sarukkai, and Pankaj Mehra. Automated performance prediction of message-passing parallel programs. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/450_SSAR/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Crandall:1995:IOC

[CACR95] Phyllis E. Crandall, Ruth A.

Aydt, Andrew A. Chien, and Daniel A. Reed. Input/output characteristics of scalable parallel applications. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/613_DREE/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Chilimbi:1995:SWA

[CBEL95] Trishul M. Chilimbi, Thomas Ball, Stephen G. Eick, and James R. Larus. Storm Watch: a tool for visualizing memory system protocols. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/505_

TCHI/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Christoph:1995:UMD

[CJN^{+}95] Gary G. Christoph, Kathleen A. Jackson, Michael C. Neuman, Christine L. B. Siciliano, Dennis D. Simmonds, Cathy A. Stallings, and Joseph L. Thompson. UNICORN: Misuse detection for UNICOS(TM). In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/714_GGC/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Clay:1995:DCP

[CS95] Robert L. Clay and Peter A.

Steenkiste. Distributing a chemical process optimization application over a gigabit network. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/592_RLCL/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Clark:1995:OLP

[CSWM95] Terry W. Clark, L. Ridgway Scott, Stanislaw Wloked, and J. Andrew McCammon. I/O limitations in parallel molecular dynamics. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/524_

TCLA/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Coppola:1995:DCS

[CT95] Ralph K. Coppola and Eva Erdosne Toth. Developing computational science curricula: The EarthVision experience. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/595_ETOT/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Curtis:1995:ISN

[Cur95] Pavel Curtis. Invited speaker: Network places: Concepts, experiences, and plans. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/MISC/INVSPEAK.

HTM#Curtis. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Das:1995:IAF

[DHSK95] Raja Das, Paul Havlak, Joel Saltz, and Ken Kennedy. Index array flattening through program transformation. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/546_HAVL/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Donegan:1995:ISH

[Don95a] Jack Donegan. Invited speaker:

How many miles per gallon does your computer get? In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/MISC/INVSPEAK.

HTM#Donegan. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Donegan:1995:MP

[Don95b] Jack Donegan (moderator).

Mayors’ panel. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/MISC/
PANELS.HTM#Mayors. These pro-

ceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Demmel:1995:PPG

[DS95] James Demmel and Sharon Smith. Performance of a parallel global atmospheric chemical tracer model. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/608_

SHSM/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Erlichson:1995:BCS

[ENSO95] Andrew Erlichson, Basem A. Nayfeh, Jaswinder P. Singh, and Kunle Olukotun. The benefits of clustering in shared address space multiprocessors: An applications-driven investigation. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/532_

AERL/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Fallah-Adl:1995:EAA

[FAJL^{+}95] Hassan Fallah-Adl, Joseph JaJa, Shunlin Liang, Yoram J. Kauman, and John Townshend. Efficient algorithms for atmospheric correction of remotely sensed data. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/511_

HFAD/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Fernandez:1995:MCA

[FM95] Jose Roberto Fernandez and Matt W. Mutka. Model and call admission control for distributed applications with correlated bursty traffic. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/502_

JRFE/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Gannon:1995:OOP

[Gan95a] Dennis Gannon. Object oriented parallel programming. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/902_DG/SC95.HTM.

These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Gannon:1995:WOO

[Gan95b] Dennis Gannon. Workshop: Object oriented parallel programming. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/902_

DG/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Garcia:1995:NAT

[Gar95] Jordi Garcia. A novel approach towards automatic data distribution. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/437_

JGAR/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Gupta:1995:PAF

[GK95] Anshul Gupta and Vipin Kumar. Parallel algorithms for forward and back substitution in direct solution of sparse linear systems. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/030_

AGUP/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Grama:1995:PMP

[GKS95] Ananth Grama, Vipin Kumar, and Ahmed Sameh. Parallel matrix-vector product using approximate hierarchical methods. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/032_

AGRA/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Gu:1995:SAD

[GLL95] Junjie Gu, Zhiyuan Li, and Gyungho Lee. Symbolic array dataflow analysis for array privatization and program parallelization. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/436_

JGU/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Gupta:1995:HCI

[GMS^{+}95] Manish Gupta, Sam Midkiff, Edith Schonberg, Ven Seshadri, David Shields, Ko-Yang Wang, Wai-Mee Ching, and Ton Ngo. An HPF compiler for the IBM SP2. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/417_SAMM/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Goodman:1995:ISE

[Goo95] Seymour Goodman. Invited speaker: An examination of high-performance computing export control policy in the 1990s. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/MISC/INVSPEAK.

HTM#Goodman. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Gannon:1995:WIS

[GSS95] Dennis Gannon (Moderator), Larry Smarr (panellist), and Vince Schuster (panellist). Where is the supercomputer software revolution? In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/MISC/
PANELS.HTM#Gannon. These pro-

ceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Hall:1995:DCG

[HAM^{+}95] Mary W. Hall, Saman P. Amarasinghe, Brian R. Murphy, Shih-Wei Liao, and Monica S. Lam. Detecting coarse-grain parallelism using an interprocedural parallelizing compiler. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/622_MHAL/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Huang:1995:MVT

[HHM95] Y. Huang, C. C. Huang, and P. K. McKinley. Multicast virtual topologies for collective communication in MPCs and ATM clusters. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/435_

HUAN/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Hendrickson:1995:MAP

[HL95] Bruce Hendrickson and Robert Leland. A multi-level algorithm for partitioning graphs. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/509_BHEN/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Hodur:1995:ISN

[Hod95] Richard M. Hodur. Invited speaker: Numerical weather prediction and the America’s Cup. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/MISC/INVSPEAK.

HTM#Hodur. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Hao:1995:CAS

[HS95a] Ming-Hong Hao and Harold A.

Scheraga. Computational approach to the statistical mechanics of protein folding. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/004_SHER/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Harzallah:1995:PAB

[HS95b] Karim Harzallah and Kenneth C. Sevcik. Predicting application behavior in large scale shared-memory multiprocessors. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/601_

KHAR/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Hemy:1995:GOD

[HS95c] Michael Hemy and Peter Steenkiste. Gigabit I/O for distributedmemory machines: Architecture and applications. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/570_MHEM/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Iwata:1995:ISV

[Iwa95] Shuichi Iwata. Invited speaker:

Virtual engineering: Challenges into handy engineering from advanced and super technologies. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/MISC/
INVSPEAK.HTM#Iwata. These

proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Jayasimha:1995:PNC

[JHP95] D. N. Jayasimha, M. E. Hayder, and S. K. Pillay. Parallelizing Navier–Stokes computations on a variety of architectural platforms. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/713_

DNJA/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Jones:1995:PPP

[JY95] Jeff A. Jones and Katherine A. Yelick. Parallelizing the phylogeny problem. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/562_JJON/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Kim:1995:ESL

[KA95] WooYoung Kim and Gul Agha. Efficient support of location transparency in concurrent object-oriented programming languages. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/425_

WYK/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Karin:1995:PAI

[Kar95] Sidney Karin, editor. Proceedings of the 1995 ACM/IEEE Supercomputing Conference, December 3–8, 1995, San Diego Convention Center, San Diego, CA, USA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/. These

proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Kohn:1995:PSI

[KB95] Scott R. Kohn and Scott B.

Baden. A parallel software infrastructure for structured adaptive mesh methods. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/507_SKOH/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Karypis:1995:AMG

[KK95a] George Karypis and Vipin Kumar. Analysis of multilevel graph partitioning. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/

                                        proceedings/031_GKAR/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Kennedy:1995:ADL

[KK95b] Ken Kennedy and Ulrich Kremer. Automatic data layout for High Performance Fortran. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/580_UKRE/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Kumaran:1995:AFE

[KMQ95] Santhosh Kumaran, Robert N. Miller, and Michael J. Quinn. Architecture-adaptable finite element modelling: A case study using an ocean circulation simulation. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/559_

SKUM/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Kuba:1995:SEC

[KPG95] Masayuki Kuba, Constantine D.

Polychronopoulos, and Kyle Gallivan. The synergetic effect of compiler, architecture, and manual optimizations on the performance of CFD on multiprocessors. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/025_

KUBA/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Kontothanassis:1995:LRC

[KSB95] Leonidas I. Kontothanassis, Michael L. Scott, and Ricardo Bianchini. Lazy release consistency for hardwarecoherent multiprocessors. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/554_KTHA/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Labarta:1995:NAT

[Lab95] Jesus Labarta. A novel approach towards automatic data distribution. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/437_

JGAR/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Lu:1995:MPV

[LDCZ95] Honghui Lu, Sandhya Dwarkadas, Alan L. Cox, and Willy Zwaenepoel. Message passing versus distributed shared memory on networks of workstations. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/586_

HLU/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Lou:1995:PIF

[LF95] John Z. Lou and Robert D.

Ferraro. A parallel incompressible flow solver package with a parallel multigrid elliptic kernel. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/018_

LOU/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Lilly:1995:UCA

[Lil95] H. Albert Lilly. The use of cellular automata in the classroom. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/482_ALIL/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Lumetta:1995:TMP

[LKC95] Steven S. Lumetta, Arvind Krishnamurthy, and David E. Culler. Towards modeling the performance of a fast connected components algorithm on parallel machines. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.

                                        org/sc95/proceedings/465_

SLUM/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Lund:1995:EAH

[LM^{+}95] Craig Lund (moderator), José Muñoz (panelist), et al. Embedded applications for high performance computing. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/804_CL/SC95.HTM.

These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Mills:1995:LTK

[MFC^{+}95] Kim Mills, Geoffrey Fox, Paul Coddington, Barbara Mihalas, Marek Podgorny, Barbara Shelly, and Steven Bossert. The living textbook and the K–12 classroom of the future. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/620_KMIL/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Milosevich:1995:IST

[Mil95] Sam Milosevich. Invited speaker:

Thriving on information anxiety: A survival guide to the knowledge-value revolution. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/MISC/INVSPEAK.

HTM#Milos. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Miller:1995:PPS

[MPP^{+}95] Craig Miller, David G. Payne, Thanh N. Phung, Herb Siegel, and Roy Williams. Parallel processing of spaceborne imaging radar data. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/012_

PAYN/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Messina:1995:WSS

[MPS95] Paul Messina, James C. T. Pool, and Thomas Sterling. Workshop: System software and tools for high-performance computing environments. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/MISC/
                                        WORKSHOP.HTM#Tools. These

proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Makino:1995:ANS

[MT95] Junichiro Makino and Makoto Taiji. Astrophysical N-body simulations on the GRAPE-4 special-purpose computer. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/721_JMAK/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Natarajan:1995:CSP

[NK95] Ramesh Natarajan and Dilip Krishnaswamy. A case study in parallel scientific computing: The boundary element method on a distributed-memory multicomputer. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/438_

KRIS/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Nassersharif:1995:HCA

[NMkL^{+}95] Bahram Nassersharif, Richard Marciano, Sui ky Ling, Eugene Ho, and Curt Edmonds. High-performance computing approaches for using the WWW to access a large-scale environmental dataset repository. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/545_RMAR/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Pancake:1995:ISE

[Pan95] Cherri Pancake. Invited speaker:

The Emperor has no clothes:

What HPC users need to say and HPC vendors need to hear. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/MISC/
                                        INVSPEAK.HTM#Pancake. These

proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Parker:1995:SAS

[PJ95] Steven G. Parker and Christopher R. Johnson. SCIRun: A scientific programming environment for computational steering. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/499_

SPAR/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Plevyak:1995:HEM

[PKZC95] John Plevyak, Vijay Karamcheti, Xingbin Zhang, and Andrew A. Chien. A hybrid execution model for fine-grained languages on distributed memory multicomputers. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/590_JPLE/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Pakin:1995:HPM

[PLC95] Scott Pakin, Mario Lauria, and Andrew Chien. High performance messaging on workstations: Illinois Fast Messages (FM) for Myrinet. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/567_SPAK/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Pool:1995:WRI

[Poo95] James C. T. Pool. Workshop: Research issues in scalable I/O. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/MISC/
WORKSHOP.HTM#Scalable. These

proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Porto:1995:PSC

[Por95] Casey Porto. Pittsburgh Supercomputing Center High School Initiative in Computational Science Report on Findings School Years: 1991–92, 1992–93, 1993– 94. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/568_CPOR/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Rinard:1995:COP

[Rin95] Martin C. Rinard. Communication optimizations for parallel computing using data access information. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/539_

MRIN/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Ramachandran:1995:AME

[RSS^{+}95] Umakishore Ramachandran, Gautam Shah, Anand Sivasubramaniam, Aman Singla, and Ivan Yanasak. Architectural mechanisms for explicit communication in shared memory multiprocessors. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/561_

GSHA/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Snoad:1995:MGC

[SB95] Nigel Snoad and Terry Bossomaier. MONSTER — the ghost in the Connection Machine: Modularity of neural systems in theoretical evolutionary research. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/531_

NSNO/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Smith:1995:TCF

[SBW95] Norris Parker Smith (moderator), Forest Baskett (panelist), and Irving Wladawski (panelist). Are tereflops commercial flops? In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/MISC/
                                        PANELS.HTM#Smith. These pro-

ceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Schill:1995:ISJ

[Sch95] John Schill. Invited speaker: Joint Task Force Advanced Technology Demonstration (JTF ATD). In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/MISC/
                                        INVSPEAK.HTM#Schill. These

proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Seamons:1995:SCO

[SCJ^{+}95] K. E. Seamons, Y. Chen, P. Jones, J. Jozwiak, and M. Winslet. Server-directed collective I/O in Panda. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/520_SEAM/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

John:1995:WGN

[SD95] Wallace B. St. John and David H. DuBois. Wide-area gigabit networking: Los Alamos HIPPI-SONET gateway. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/710_WSTJ/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Skordos:1995:PPS

[Sko95] Panayotis Skordos. Price and performance of simulating wind instruments. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/722_

PSKO/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Smith:1995:WHA

[Smi95] Brian T. Smith. Workshop: HPF: A user’s perspective. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/MISC/WORKSHOP.

HTM#HPF. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Siegell:1995:CAG

[SS95] Bruce S. Siegell and Peter A. Steenkiste. Controlling application grain size on a network of workstations. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/591_

BSIE/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Sterling:1995:PEC

[SSM^{+}95] Thomas Sterling, Daniel Savarese, Peter MacNeice, Kevin Olson, Clark Mobarry, Bruce Fryxell, and Phillip Merkey. A performance evaluation of the Convex SPP-1000 scalable shared memory parallel computer. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/709_PMER/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Stewart:1995:HUC

[Ste95] Kris Stewart. HPC undergraduate curriculum development at SDSU using SDSC resources. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/704_

KSTE/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Bruggencate:1995:PIP

[tBC95] Monika ten Bruggencate and Suresh Chalasani. Parallel implementations of the power system transient stability problem on clusters of workstations. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/600_MBRU/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Teller:1995:MRT

[TD95] Patricia J. Teller and Ted Dunning. Mobile robots teach machine-level programming. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/433_TELL/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Eyck:1995:SMI

[TMRP95] Lynn F. Ten Eyck, Jeffrey Mandell, Victoria A. Roberts, and Michael E. Pique. Surveying molecular interactions with DOT. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/636_

LTEN/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Topham:1995:COD

[TRM^{+}95] Nigel Topham, Alasdair Rawsthorne, Callum McLean, Muriel Mewissen, and Peter Bird. Compiling and optimizing for decoupled architectures. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/471_TOPH/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Villacis:1995:WIP

[VG95] Juan Villacis and Dennis Gannon. A Web interface to parallel program source code archetypes. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/564_

JVIL/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Williams:1995:SFU

[WB95] Alan Williams and Kevin Burrage. Surface fitting using GCV smoothing splines on supercomputers. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/494_

ABW/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Waheed:1995:SAI

[WR95] Abdul Waheed and Diane T. Rover. A structured approach to instrumentation system development and evaluation. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

                                        //www.supercomp.org/sc95/
                                        proceedings/609_WAHE/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Wray:1995:WEA

[Wra95] Francis Wray. Workshop: EU-ROPORT activities. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/MISC/WORKSHOP.

HTM#EUROPORT. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Wu:1995:HIS

[WS95] Min-You Wu and Wei Shu.

High-performance incremental scheduling on massively parallel computers — A global approach. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/477_

WSHU/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Wu:1995:LES

[WSL95] Xiaohua Wu, Kyle D. Squires, and Thomas S. Lund. Large eddy simulation of a spatiallydeveloping boundary layer. In Karin [Kar95], page ?? ISBN 0-89791-816-9. URL http:

//www.supercomp.org/sc95/
proceedings/428_SQUI/SC95.

HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Wulf:1995:NSR

[Wul95] William A. Wulf. And now for some *really* super computing. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/MISC/

KEYNOTE.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Yoshida:1995:QCS

[YNF^{+}95] M. Yoshida, A. Nakamura, M. Fukuda, T. Nakamura, and S. Hioki. Quantum chromodynamics simulation on NWT. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

                                        URL http://www.supercomp.
                                        org/sc95/proceedings/720_

MF/SC95.HTM. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Ziebarth:1995:CNA

[Zie95a] John Ziebarth (moderator).

Community networking I — applications. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/MISC/
PANELS.HTM#CNetI. These pro-

ceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Ziebarth:1995:CNI

[Zie95b] John Ziebarth (moderator).

Community networking II — technology. In Karin [Kar95], page ?? ISBN 0-89791-816-9.

URL http://www.supercomp.
org/sc95/proceedings/MISC/
PANELS.HTM#CNetII. These pro-

ceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).